Mike's ISF is 40 and his insulin to CHO ratio is 1:12. What is the correct pre-meal rapid acting insulin dose for Mike based on the following. Round to nearest whole number.

Pre meal - BG 102 mg/dI

Meal - 4 ounces grilled chicken

1.5 cups rice

1 teaspoon butter
Answer #1

Answer : Pre meal sugar for Mike = 102mg/dl

Insulin ratio of 1:12 indicates 1 unit of insulin covers 12 grams of carbohydrate.

carbohydrate in Breakfast :

4 ounces of grilled chicken= 0.1 grams CHO

1.5 cup rice=68grams CHO

1 teaspoon butter= 0 grams CHO

Total amount of carbohydrate consumed in breakfast = 0.1+ 68 + 0= 68.1 grams

Amount of insulin that will cover 68.1 grams of carbohudrate = 1/2 x 68.1= 5.66Unit

Round up to the nearest whole number= 6 unit